# Comparing Neighbors Together Makes it Easy: Jointly Comparing Multiple Candidates for Efficient and Effective Retrieval

### v1

@@ 3.2 Inference
CMC is capable of pre-computing and storing candidate embeddings offline, unlike cross-encoders Logeswaran et al. (2019); Wu et al. (2020) (Figure 1). CMC processes the candidates in the collection (e.g. knowledge base) and stores a single embedding for each. As demonstrated in Section 4.4, offline indexing significantly reduces the inference time compared to that of cross-encoders. Furthermore, it allows the runtime performance of CMC to be comparable to that of bi-encoders. While reducing time complexity, the space requirement for our model is less than $1\%$ of that required by Sum-of-Max and Deformer, as these models necessitate every token embedding for scoring, whereas CMC requires only a single vector embedding per candidate.
Parallel Computation
Given multiple GPUs, the end-to-end runtime for retrieving and reranking with CMC  can be comparable to that of bi-encoders retrieval. This is achieved through parallel computation of query encoders at bi-encoders and CMC (Figure 2). Consequently, the additional latency for running CMC is limited to the execution of a few encoder stacks within the self-attention layer.
3.3 CMC as a Virtually Enhanced Retriever
The parallel computation of CMC (Section 3.2) offers an experience of virtually enhanced retriever, requiring only a slight increase in computing resources (virtual) while significantly improving recall at various K (↑R@K) (enhanced). This process begins with initial retrievers such as bi-encoders, which retrieve a broad set of candidates. CMC then narrows this set down to a more manageable number for cross-encoders (e.g., 64 or fewer). Since CMC can rerank candidates from the initial retriever with only a marginal increase in latency, the runtime for bi-encoders is comparable to that of a virtually enhanced retriever running bi-encoders and CMC in parallel. Consequently, the improved recall contributes to the performance increase of the final stage reranker (e.g., cross-encoders) at a marginal cost, optimizing overall performance with comparable efficiency (Section 4.3).
3.4 CMC as a Reranker
The obvious application of CMC is the final stage reranker whose goal is to increase R@1 as well as R@K. Effective reranking is achieved by contextualizing the query and the candidates concerning each other while maintaining efficiency by utilizing a single vector embedding for each. In the training stage, CMC is typically fed with 64 candidates per query. Surprisingly, this training setup remains effective for a varying number of candidates at the inference stage. For example, even though CMC was trained with 64 candidates on the MS MARCO passage ranking dataset, it shows effectiveness when handling up to 1,000 candidates (Section §4.4). This evidence not only emphasizes the scalability of CMC but also its robustness in processing a diverse distribution of candidates.
@@ 4 Experiments
4.1 Dataset
To assess the empirical impact of CMC , we conduct our experiments on entity linking, passage ranking, and dialogue ranking tasks, where the retrieve-and-rerank approach is commonly employed. For entity linking, we utilize datasets linked to the Wikipedia knowledge base (AIDA-CoNLL Hoffart et al. (2011), WNED-CWEB Guo and Barbosa (2018), and MSNBC Cucerzan (2007)), as well as a ZEro-SHot Entity Linking dataset (ZeSHEL; Logeswaran et al. (2019)) based on the Wikia333now Fandom: https://www.fandom.com knowledge base. The candidates are retrieved from bi-encoders fine-tuned for each knowledge base Wu et al. (2020); Yadav et al. (2022). For passage ranking, we conduct an experiment on MS MARCO with 1K candidates from BM25 as an initial retriever following Bajaj et al. (2016). For dialogue ranking tasks, we test our model on DSTC7 challenge (Track 1) Yoshino et al. (2019), where conversations are extracted from technical support chat. For more details, see Appendix §C.
@@ 4.1 Dataset
Single- BM25 25.92 44.86 52.05 58.21 63.76 69.13
View Bi-encoder (BE) 52.94 64.51 71.94 81.52 84.98 87.95 568.9 0.2
Poly-encoder 40.04 60.16 67.23 72.16 76.49 80.18 581.0 0.2
Arbo-EL 50.31 68.34 74.26 78.4 82.02 85.11 - -
GER 42.86 66.48 73.00 78.11 81.11 85.65 - -
Sum-of-max 27.11 64.11 73.21 79.59 84.14 87.95 6393.0 25.7
BE + CMC(Ours) 59.08 77.58 82.91 86.32 89.33 91.51 607.2 0.4
Multi- MuVER 43.49 68.78 75.87 77.65 85.86 89.52 - -
View MVD 52.51 73.43 79.74 84.35 88.17 91.55 - -
MVD + CMC(Ours) 59.04 77.77 83.06 86.71 89.91 92.36 - -
Table 1: Retrieval performance over ZeSHEL dataset. The best result is denoted in bold and the second-best result is underlined for each method. BE is bi-encoders from Yadav et al. (2022) which is used for CMC. indicates our implementation as recall@k for all k are not provided in previous work555recall@64 of Poly-encoder and Sum-of-max from Zhang and Stratos (2021) is reported as 84.34 and 89.62, respectively.. Our result is conducted over top-512 candidates from the initial retriever and averaged over experiments with five random seeds.
@@ 4.2 Training Details
All the training details for CMC and other baselines are the same, following prior studies Wu et al. (2020); Zhang and Stratos (2021); Gao and Callan (2022); Yang et al. (2023). CMC and other baselines are trained using the AdamW optimizer with a 10% warmup step. Also, we investigate different sentence encoders for CMC and late interaction models: (1) vanilla BERT and (2) BERT-based sentence encoders fine-tuned on each dataset. After training, we select the best result among them. [^6] If more favorable results are found from prior works over the same candidates, we source the results from these works. For ZeSHEL, training CMC and other low-latency baselines for one epoch on a single NVIDIA A100 GPU takes approximately 4 hours. Every training detail for each dataset is outlined in Appendix §D.
4.3 CMC as a Virtually Enhanced Retriever
We conduct experiments on the ZeSHEL dataset to verify the impact of CMC as a virtually enhanced retriever. We conduct experiments on how the parallel computation of CMC reranking can achieve a higher recall@k with only a marginal increase in computational cost compared to conventional bi-encoders. We also evaluate how improved recall@k can enhance the performance of cross-encoders’ reranking. To do so, we compare how well cross-encoders rerank candidates filtered only by the bi-encoders versus those further refined by CMC.
@@ 4.3 CMC as a Virtually Enhanced Retriever
To assess CMC’s performance as a retriever, we compare CMC against baselines categorized into two types: single-view and multi-view retriever.  [^7] Single-view retrievers consider only a single global view derived from the entire sentence, whereas multi-view retrievers divide candidate information into multiple local views. As an initial retriever that provides candidates for CMC, we use bi-encoders fine-tuned for ZeSHEL Yadav et al. (2022) for the single-view retriever, and MVD Liu et al. (2023) for multi-view retriever. As baselines, we select baselines such as the poly-encoder Humeau et al. (2019), Sum-of-max Zhang and Stratos (2021), Arbo-EL Agarwal et al. (2022b), and GER Wu et al. (2023) for single-view retriever, and we include MuVER Ma et al. (2021) and MVD Liu et al. (2023), recognized as the SOTA retrievers for the ZeSHEL dataset.
@@
In Table 5, adopting CMC upon single-view retriever surpasses other baseline models across all k values, demonstrating its effectiveness in enhancing the end-to-end retrieval process. Also, CMC boosts recall@64 to 91.51% with candidates from the initial retriever whose recall@64 is 87.95%, with a marginal increase in latency (+0.07x). Similarly, operating CMC with MVD shows that virtually enhanced retriever consistently improves recall performance, regardless of the initial retriever used. It leads to an increase in performance, moving from 91.55 to 92.36 at recall@64. For the effect of the number of candidates from the initial retriever, see appendix §E.2.
To assess the effect of candidates retrieved from CMC on the overall performance, we analyze the subsequent cross-encoder’s reranking performance in Table 2. We explore various combinations where candidates are selected from either bi-encoders or CMC, in groups of (8, 16, 64). These candidates from CMC are initially selected from pools of (64, 256, 512) provided by bi-encoders which are the same or larger than 64 that typical cross-encoders consider.
A key finding presented in the table is that CMC’s scalability to search over a broad set enables it to provide better candidates for cross-encoders. By filtering candidates from bi-encoders to fewer, CMC effectively narrows down the most promising candidates for cross-encoders. In Table 2, the performances of cross-encoders with 16 candidates from CMC which filters out 256 or 512 candidates (row 8-9) are slightly better than their best results with 64 bi-encoder candidates (row 3), with a significant reduction in latency, making it 1.75 times faster by presenting fewer candidates for cross-encoders. Furthermore, cross-encoders reach their best performance with candidates from CMC. When using 64 candidates from CMC (row 11), cross-encoders surpass their performance with an equal number of bi-encoder candidates (row 3) by 0.17 points in macro accuracy.
@@
Table 2: Unnormalized accuracy of cross-encoders across various candidate configurations on the ZeSHEL dataset. We underlined when the cross-encoders show superior accuracy with candidates generated by CMC compared to those from bi-encoders. The top-performing scenarios in each category are highlighted in bold. We measure the comparative latency required for running cross-encoders over 64 bi-encoder candidates (260.84ms). For your reference, the CMC runtime 2x when increasing the number of candidates by 16x (from 128 to 1048), while able to compare up to 16k candidates at once. (Appendix E.1)
@@
Intermediate- Deformer 79.6$\pm$0.8 63.6±$\pm 0.3$ 23.0 35.7 68.6 76.4 4.39x 125x
Latency Sum-of-max 80.7±$\pm 0.2$ 58.8±$\pm 1.0$ 22.8 35.4 66.9 75.5 5.20x -
- w/ offline indexing 80.7±$\pm 0.2$ 58.8±$\pm 1.0$ 22.8 35.4 66.9 75.5 1.05x 125x
@@
Table 3: Reranking Performance on four datasets with three downstream tasks: Entity Linking (Wikipedia-KB based datasets Hoffart et al. (2011); Guo and Barbosa (2018); Cucerzan (2007), ZeSHEL Logeswaran et al. (2019), Passage Ranking (MS MARCO Passage Ranking Bajaj et al. (2016), and Dialogue Ranking Gunasekara et al. (2019). The best result is denoted in bold and the second-best result is underlined. MRR stands for mean reciprocal rank and R@1 stands for recall at 1. In the entity linking datasets, the results are derived from averages across five random seeds. To show the computing resources required for the reranking process, we define reranking latency in terms of relative latency and additional memory usage compared to bi-encoders. indicates that the results are sourced from Wu et al. (2020); Yang et al. (2023); Yadav et al. (2022), respectively.
4.4 CMC as a Reranker
@@
In Table 3, baseline models are categorized into high-latency, intermediate-latency, and low-latency models. We adopt cross-encoders as our primary baseline for the high-latency approach. For the intermediate-latency models, we include Deformer and Sum-of-max, which utilize all vector embeddings to represent candidate information. For the low-latency models, we include the Bi-encoder, Poly-encoder, and Mixencoder, all of which require a single vector embedding for representation and have a serving time similar to that of the Bi-encoder. In this context, CMC is classified as a low-latency method because it requires a single embedding for the candidate and takes 1.17x serving time of the Bi-encoder.
@@
Given the critical importance of computational resources and serving time in applications, CMC can be a highly practical alternative to cross-encoders. CMC shows 11.02x speedup and comparable prediction performance. CMC outperforms the cross-encoder in the Wikipedia entity linking (+0.7 accuracy points) and DSTC7 dataset (+3.3 MRR points) and presents a competitive performance in MS MARCO and ZeSHEL dataset, achieving the second- or third-best prediction performance. This comparison in performance underscores the significant role of the self-attention layer in CMC, suggesting that it effectively replaces the token-by-token interaction in cross-encoders with a balance between accuracy and computational efficiency.
Comparison with Intermediate-latency Models
When compared with intermediate-latency models such as Deformer and Sum-of-max, CMC demonstrates its capability not just in memory efficiency but also in maintaining competitive performance. CMC mostly surpasses these models in entity linking and passage ranking tasks. Also, CMC offers significant improvements in computational speed over Deformer (1.17x vs. 4.39x) and Sum-of-max without caching (1.17x vs. 5.20x).
For Sum-of-max with caching, it poses a critical issue in memory efficiency, requiring an index size 125 times larger than that of CMC. This gap becomes critical in terms of memory efficiency and scalability when managing extensive document collections. For instance, Sum-of-max and deformer require 3.2 TB to store just 5M entity descriptions from Wikipedia with 128 tokens, while CMC requires 23GB to store the same entities. As the knowledge base and document collections continuously change and grow, the operational feasibility of these systems quickly becomes impractical. Furthermore, the need to frequently update the embeddings to integrate the latest advances from pre-trained language models adds another layer of complexity and resource demand.
This analysis suggests that CMC’s single-vector approach is not only significantly faster under practical resource constraints but also demonstrates a comparable ability to represent candidate information with less information, occasionally surpassing more complex methods.
Comparison with Low-latency Models
In comparison with low-latency models, CMC matches or surpasses the performance of other baselines like Bi-encoder, Poly-encoder, and Mixencoder across different datasets. When compared with bi-encoders, the architectural distinctions contribute to CMC’s enhanced performance across every dataset. It shows the effectiveness of the CMC’s self-attention layer over the simple dot product. When evaluated against the Poly-encoder, CMC falls short on conversational datasets, as reported by Humeau et al. (2019). However, it outperforms Poly-encoder on other datasets not covered in the original baseline paper, which shows the robustness of CMC in handling more complex tasks than dialogue ranking. Similarly, CMC outperforms MixEncoder in entity linking and passage ranking.

### v2

@@ 3.3 Inference
CMC can pre-compute and index the embeddings of candidates in the collection (e.g., knowledge base), unlike cross-encoders (Figure 1). This offline indexing scheme significantly reduces inference time compared to cross-encoders, making the runtime of CMC comparable to that of bi-encoders (§4.4). While reducing time complexity, CMC is highly memory-efficient requiring less than $1\%$ of index size needed by Sum-of-Max and Deformer, which store every token embedding per candidate. This is because CMC only stores a single vector embedding per candidate.
Parallel Computation of Query Representations
The end-to-end runtime for retrieving and reranking with CMC can be comparable to that of bi-encoder retrieval. The runtime can be further improved by parallelizing query encoders in both bi-encoder and CMC (Figure 2). Ideally, the additional latency for running CMC is limited to the execution of a few self-attention layers.
CMC as the Seamless Intermediate Reranker
CMC can serve as a seamless intermediate reranker that maintains the latency-wise user experience while providing improved retrieval performance when combined with a bi-encoder. Thanks to the parallel computation discussed above, plugging in CMC after bi-encoders should minimally impact retrieval latency compared to just using the bi-encoder. The process starts with the first-stage retrievers, such as bi-encoders, retrieving a broad set of candidates. CMC then narrows this set down to fewer, higher-quality candidates with a more manageable number (e.g., 64 or fewer) for the reranker. Since CMC, the seamless intermediate reranker, filters candidates from the first-stage retriever with negligible additional latency, its runtime is comparable to that of bi-encoders. As a result, the improved candidate quality boosts the prediction accuracy of the final-stage reranker (e.g., cross-encoders) with only a marginal increase in computational cost (Figure 3; §4.3).
CMC as the Final Stage Reranker
CMC can obviously serve as the final-stage reranker to increase top-1 accuracy. Enriching contextualized representations of the query and candidates helps improve top-1 accuracy in reranking while maintaining efficiency with a single vector embedding. Notably, CMC remains effective even when the number of candidates varies during inference, despite being trained with a fixed number of candidates. For example, when trained with 64 candidates on the MS MARCO passage ranking dataset, CMC still performs effectively with up to 1K candidates. This demonstrates not only the scalability of CMC but also its robustness in processing a diverse range of candidate sets (§4.4).
@@ 4 Experiments
Single- BM25 25.9 44.9 52.1 58.2 63.8 69.1
View Bi-encoder (BE) 52.9 64.5 71.9 81.5 85.0 88.0 568.9 0.2
Arbo-EL 50.3 68.3 74.3 78.4 82.0 85.1 - -
GER 42.9 66.5 73.0 78.1 81.1 85.7 - -
Poly-encoder (Poly) 40.0±$\pm 0.7$ 60.2±$\pm 0.9$ 67.2±$\pm 0.7$ 72.2±$\pm 0.8$ 76.5±$\pm 0.8$ $\pm 0.8$ 581.0 0.2
BE + Poly 56.9±$\pm 0.8$ 74.8±$\pm 0.6$ 80.1±$\pm 0.7$ 84.2±$\pm 0.5$ 87.5±$\pm 0.4$ 90.2±$\pm 0.3$ 574.6 0.4
Sum-of-max (SOM) 27.1±$\pm 1.8$ 64.1±$\pm 1.4$ 73.2±$\pm 0.9$ 79.6±$\pm 0.7$ 84.1±$\pm 0.4$ 88.0±$\pm 0.4$ 6393.0 25.7
BE + SOM 58.5±$\pm 1.0$ 76.2±$\pm 1.1$ 81.6±$\pm 1.0$ 85.8±$\pm 0.9$ 88.9±$\pm 0.7$ 91.4±$\pm 0.6$ 2958.3 0.2
- w/ offline indexing 597.3 25.9
BE + CMC(Ours) 59.1±$\pm 0.3$ 77.6±$\pm 0.3$ 82.9±$\pm 0.1$ 86.3±$\pm 0.2$ 89.3±$\pm 0.2$ 91.5±$\pm 0.1$ 607.2 0.4
Multi- MuVER 43.5 68.8 75.9 77.7 85.9 89.5 - -
View MVD 52.5 73.4 79.7 84.4 88.2 91.6 - -
MVD + CMC(Ours) 59.0 77.8 83.1 86.7 89.9 92.4 - -
Table 1: Retrieval performance over ZeSHEL dataset. The best and second-best results are denoted in bold and underlined. BE is bi-encoder from Yadav et al. (2022) which is used for CMC. indicates our implementation as recall@k for all k are not provided in previous work777recall@64 of Poly-encoder and Sum-of-max from Zhang and Stratos (2021) is reported as 84.34 and 89.62, respectively.. results on BE + Reranker (e.g., BE+CMC) are conducted over the top 512 candidates from the first-stage retriever and averaged over experiments with 5 random seeds.
4.1 Dataset
To evaluate the robustness of CMC, we conduct experiments on various ranking tasks where the retrieve-and-rerank approach is commonly used. For entity linking, we utilize datasets linked to the Wikipedia knowledge base (AIDA-CoNLL Hoffart et al. (2011), WNED-CWEB Guo and Barbosa (2018), and MSNBC Cucerzan (2007)), as well as a ZEro-SHot Entity Linking dataset (ZeSHEL; Logeswaran et al. (2019)) based on the Wikia888now Fandom: https://www.fandom.com knowledge base. The candidates are retrieved from bi-encoders fine-tuned for each knowledge base Wu et al. (2020); Yadav et al. (2022). For passage ranking, we conduct an experiment on MS MARCO with 1K candidates from BM25 as the first-stage retriever following Bajaj et al. (2016). For dialogue ranking tasks, we test our model on DSTC7 challenge (Track 1) Yoshino et al. (2019), where candidates are officially provided. The primary metric used is recall@k, as datasets typically have only one answer or rarely a few answers per query. Further details are presented in §B.
@@ 4.2 Training Details
CMC and other baselines are trained under the same training strategies. All models use the same loss function and negative sampling (§3.2) with the AdamW optimizer and a 10% linear warmup scheduler. Also, we examine diverse sentence encoder initialization for CMC and late interaction models, including vanilla BERT and BERT-based models fine-tuned on in- and out-of-domain datasets. After training, we select the best results for each model. [^9] If more favorable results are found in prior works over the same candidates, we use those results. For ZeSHEL, training CMC and other low-latency baselines for one epoch on an NVIDIA A100 GPU takes about 4 hours. The training details for each dataset are in §C, and the ablation studies for diverse training strategies are presented in §4.5 and §D.5.
4.3 CMC as the Seamless Intermediate Reranker
We conduct two experiments on the ZeSHEL dataset to verify the impact of CMC as the seamless intermediate retriever (BE+CMC+CE). We examine whether the introduction of CMC can improve retrieval performance with negligible overhead as promised. In the first experiment, we compare the performance and speed of CMC plugged in with bi-encoders (BE+CMC) with other retrieval pipelines. Remarkably, even when other rerankers are plugged in with the same bi-encoder, CMC still achieves the highest Recall@k (Table 7) at a marginal latency increase. In the second experiment, we assess how a more confident set of candidates retrieved by BE+CMC contributes to improving end-to-end (BE+CMC+CE) accuracy compared to solely using bi-encoders (Figure 3).
@@ 4.3 CMC as the Seamless Intermediate Reranker
To assess CMC’s effectiveness in enhancing retrieval, we evaluate BE+CMC on 512 bi-encoder retrieved candidates and compare it to baselines categorized into two types: single- and multi-view retrievers. [^10] Single-view retrievers consider only a single global view derived from the entire sentence, whereas multi-view retrievers divide candidate information into multiple local views. We use bi-encoders Yadav et al. (2022) and MVD Liu et al. (2023) as the first-stage retrievers for the single-view and multi-view settings, respectively. For the baselines, we select the state-of-the-art retrievers for the ZeSHEL dataset. For single-view retrievers, we select the poly-encoder Humeau et al. (2019), Sum-of-max Zhang and Stratos (2021), Arbo-EL Agarwal et al. (2022b), and GER Wu et al. (2023). Among these, Arbo-EL and GER utilize graph information, unlike CMC and other baselines. For multi-view retrievers, we include MuVER Ma et al. (2021) and MVD Liu et al. (2023).
@@
In Table 7, plugging in CMC with a single-view retriever outperforms baselines across all $k$, demonstrating its effectiveness in the end-to-end retrieval process. With a marginal increase in latency (+0.07x), CMC boosts recall@64 to 91.51% on the candidates from the first-stage retriever, which has a recall@64 of 87.95%. Especially, the recall of Poly-encoder and Sum-of-max lags behind CMC even when they are plugged in with the same bi-encoders (BE+Poly & BE+SOM). Sum-of-max, which closely follows CMC, requires a tremendous index (60x of CMC) to achieve comparable latency to CMC. To show that CMC seamlessly enhances any retriever type, we examine the increase in recall of CMC upon a multi-view retriever (MVD+CMC). The results show that CMC consistently improves recall performance, moving from 91.55% to 92.36% at recall@64. This demonstrates CMC’s general capability to enhance recall performance, regardless of the first-stage retriever. For the effect of the number of candidates from the first-stage retriever, see §D.2.
Figure 3: Illustration of candidate retrieval for cross-encoders (CE). Suppose cross-encoders can process up to M candidates due to limited scalability. (a) In bi-encoder (BE) retrieval, the BE-CE framework takes M candidates and risks missing the gold candidates due to inaccurate bi-encoders, causing the entire system to suffer from error propagation from the retriever and fail to get the correct candidate. (b) When CMC is introduced as the seamless intermediate reranker (BE-CMC-CE), CMC can consider a significantly larger pool (K) of BE candidates. This allows CMC to provide much fewer K’ (K>M>K’) and higher-quality candidates to the CE while increasing the chance to include the positive candidate.
We question whether BE+CMC can reduce the latency of the overall retrieval and reranking process while maintaining the overall accuracy (Figure 3). In essence, if we can have fewer but higher quality candidates, end-to-end accuracy can be improved while CE forward passes are called fewer times with a reduced set of candidates. To examine the quality of candidates from the seamless intermediate reranker CMC, we report the final reranking accuracy of cross-encoders when candidates are retrieved by BE+CMC and compare it to conventional BE retrieval (Table 12).
Table 12 shows that cross-encoders outperform conventional bi-encoders, even with fewer candidates retrieved by CMC. Cross-encoders with 16 candidates from CMC are 1.75x faster and achieve slightly better accuracy compared to using 64 bi-encoder candidates (line 3 vs. 8-9). Furthermore, cross-encoders reach the best accuracy with 64 candidates from CMC, surpassing the accuracy obtained with the same number of bi-encoder candidates, with only a marginal increase in latency (line 3 vs. 11).
@@
Table 2: Unnormalized accuracy [^12] The unnormalized accuracy of the reranker in ZeSHEL is defined as the accuracy computed on the entire test set. In contrast, the normalized accuracy is evaluated on the subset of test instances for which the gold entity is among the top-k candidates retrieved by the initial retriever. For example, if the retriever correctly identifies candidates for three out of five instances and the reranker identifies one correct candidate, unnormalized accuracy is 1/5 = 20%, and normalized accuracy is 1/3 = 33%. of cross-encoders across various candidate configurations on the ZeSHEL dataset. We underlined when the cross-encoders show superior accuracy with candidates filtered by CMC compared to those from bi-encoders. The top-performing scenarios in each category are highlighted in bold. We measure the comparative latency required for running cross-encoders over 64 bi-encoder candidates (260.84ms). For your reference, the CMC runtime 2x when increasing the number of candidates by 16x (from 128 to 1048), while able to compare up to 16k candidates at once. (§D.1)
@@
Medium-latency Deformer 79.6$\pm$0.8 63.6±$\pm 0.3$ 23.0 35.7 68.6 76.4 4.39x 125x
Sum-of-max 80.7±$\pm 0.2$ 58.8±$\pm 1.0$ 22.8 35.4 66.9 75.5 5.20x -
- w/ offline indexing 1.05x 125x
@@
Table 3: Reranking Performance on four datasets with three downstream tasks: Entity Linking (Wikipedia-KB based datasets Hoffart et al. (2011); Guo and Barbosa (2018); Cucerzan (2007), ZeSHEL Logeswaran et al. (2019), Passage Ranking (MS MARCO Passage Ranking Bajaj et al. (2016), and Dialogue Ranking Gunasekara et al. (2019). The best result is denoted in bold and the second-best result is underlined. MRR stands for mean reciprocal rank. In the entity linking datasets, the results are averaged across five random seeds. To show the computing resources required for the reranking process, we define reranking latency in terms of relative latency and additional memory usage compared to bi-encoders. indicates that more favorable results are sourced from Wu et al. (2020); Yang et al. (2023); Yadav et al. (2022), respectively.
4.4 CMC as the Final Stage Reranker
@@
Baselines are categorized into high-, medium-, and low-latency models. We adopt cross-encoders as our primary baseline for the high-latency model. For the medium-latency models, we include Deformer and Sum-of-max, which utilize all token embeddings to represent candidate information. For the low-latency models, we include the Bi-encoder, Poly-encoder, and Mixencoder, all of which require a single vector embedding for representation and have a serving time similar to that of the Bi-encoder. In this context, CMC is classified as a low-latency method because it requires a single embedding for the candidate and takes 1.17x serving time of the Bi-encoder.
Comparison with Low-latency Models
CMC is highly effective across diverse datasets, outperforming or being comparable to other low-latency baselines. Notably, CMC surpasses bi-encoders on every dataset with only a marginal increase in latency. This indicates that replacing simple dot products with self-attention layers across multiple candidates can enhance reranking performance, likely by taking advantage of the relational dynamics among the candidates. Evaluated against the Poly-encoder and MixEncoder, CMC demonstrates superior prediction capability in tasks like passage ranking and entity linking, which require advanced reading comprehension capability.
Comparison with Medium-latency Models
When compared with Medium-latency models such as Deformer and Sum-of-max, CMC demonstrates its capability not only in memory efficiency but also in maintaining strong performance. CMC mostly surpasses these models in entity linking and passage ranking tasks. Also, CMC offers significant improvements in speed over Deformer (1.17x vs. 4.39x) and Sum-of-max without caching (1.17x vs. 5.20x). For Sum-of-max with caching, it requires a huge memory index size (125x) to accomplish a similar latency to CMC. If a 125x index size is not available in practice, the speed becomes impractical introducing scalability limitations. This analysis implies that CMC’s single-vector approach is significantly faster and more memory efficient, while still demonstrating a comparable ability to represent candidate information with fewer tokens, often surpassing more complex methods.
@@
Given the importance of computational resources and serving time in applications, CMC is a practical alternative to cross-encoders, with 11.02x speedup and comparable reranking accuracy. CMC outperforms the cross-encoder in the Wikipedia entity linking (+0.7p accuracy) and DSTC7 dialogue ranking (+3.3p MRR). Also, CMC presents a competitive result in MS MARCO and ZeSHEL dataset, achieving the second- or third-best in prediction. This comparison suggests that the self-attention layer in CMC effectively substitutes for the token-by-token interaction in cross-encoders while enhancing the computational efficiency of the reranking process.
In summary, to achieve the best accuracy, we recommend the 3-stage retrieval pipeline of bi-encoders + CMC + cross-encoders (BE-CMC-CE) that is both more accurate and substantially faster than the widely adopted bi-encoder + cross-encoder (BE-CE), as shown in Table 12 and §4.3. If there’s a time constraint, using CMC as the final reranker can be a good option since inferring with 16 candidates using a cross-encoder takes approximately the same amount of time as comparing around 10K candidates with CMC (Figure 4)